OpenStudy (reneeshumpert):

y + 3 = 2x y + 4 = 3x If you use the comparison method to solve the given system, what is the resulting equation? y + 3 = y + 4 2x - 3 = 3x - 4 2x + 3 = 3x + 4 Okay I struggled with this section once before, I'm not to good at the method, I just need someone to walk me through. Please do not give me the answer I'm actually trying to learn thank you

OpenStudy (anonymous):

y + 3 = 2x or, y=2x-3 y + 4 = 3x or, y=3x-4 so by comparing, 2x-3=3x-4

OpenStudy (reneeshumpert):

Thank you , because I was confused.

OpenStudy (anonymous):

k